OpenStudy (anonymous):

A=1/2bh area of a triangle can this be turned to a binomial to find the area?

OpenStudy (jdoe0001):

hmmm why do you need it as binomial anyway?

OpenStudy (anonymous):

Instructions say "The measurements of your shape should be binomials" the word "should" is telling me that this cant be turned to a binomial since its a trinomial. I believe I am just over analyzing.

OpenStudy (jdoe0001):

sounds to me more like it is highly likely it iS a binomial what's the whole enchilada?

OpenStudy (jdoe0001):

as a matter of binomial or else one could say, any number or variable plus 0 is equal to itself so 25 = 25+0 = 25 112345 = 112345 + 0 = 112345 whatever = whatever + 0 = whatever and those are binomials

OpenStudy (jdoe0001):

by the same token you could just make them trinomials or whatever else 2 = 2 + 0 + 0 = 2 and so on

OpenStudy (anonymous):

So to turn this to a binomial it would be 1/2bh+0 which is a binomial- this being my intial equation would be my binomial b=3 h=5 after substituting it would be 1/2*3*5 which is trinomial